LSC-Virgo Burst Analysis Working Group

Navigation

Burst Group Home
Wiki
ExtTrig Home
CBC Group Home
LSC, LIGO, Virgo

Documents

'How-to' docs
Agendas/minutes
[2008, 2007, earlier]
Talks/Posters [pre-wiki]
Papers
Paper plans
White papers
ViewCVS

Investigations

Analysis projects
Old notebook [General, S2, S3, S4, S5]
Virgo workarea
External collabs

Review

Main review page
Telecons

Review Committee Meeting Monday February 9 2009 08:00 Pacific / 11:00 Eastern

Minutes: Monday February 9 2009 08:00 Pacific / 11:00 Eastern

Agenda and Contact Info

Agenda

  1. Announcements
  2. Status of S5y1 all-sky search papers
    • Low-frequency search [ PDF ]
  3. Status of S5y2/VSR1 all-sky search paper
  4. A.O.B.

Contact Info

InterCall telecon service:
Phone: 1-866-380-5536, participant code: 435 672 9587 #
International callers ++1-816-249-4731 with same code
International access numbers [ HTML ]
Handy participant command: Press *6 ["*M"] to mute, #6 to unmute

Minutes

  1. Announcements
  2. Status of S5y1 all-sky search papers
    • High-frequency search.
      • Peter Saulson took issue with idea of being able to detect anything in H1/H2 datastream at all. Without null stream.
      • Null stream would have been used if there was an outlier. Discussion of how outliers would have been treated is not really clear.
      • Peter Shawhan has a potential change that might ease concerns.
      • Hopefully can circulate to LSC later this week.
      • HF calibration expected by end of Feb.
    • Low-frequency search [ PDF ]
      • We're now of the same mind about how to finalize the low-frequency paper.
      • cWB section needs work.
      • Jolien concerned about Eq. (B1) of the Q-transform... is this the windowing function?
      • (Not clear if most recent vesion is available from webpage though.)
      • Hopefully can circulate to LSC soonish, next week?
      • Who has checked the calculation of the final UL. Who has checked the livetimes etc.? This needs to be checked.
  3. Status of S5y2/VSR1 all-sky search paper
    • Feb 4 taget date has slipped by a week. Sergey is still trying to collect material. Abstract has agreement from a good fraction of the group. Discussion section also being drafted. This is needed for conclusion of blind injection challenge.
    • Still using V3 calibration--even for HF--shown to be robust. Will need to do some comparisons between V3 and V4 calibration. May not be so happy with coherent searches at high frequency using V3 because of the phase uncertainty.
  4. A.O.B.
    • Position reconstruction review.
      • Visiting scientist at MIT, Xiaoge Wang, available to do this kind of review.
      • Perhaps look at all the different methods. cWB, Omega, triangulation.
      • Need to have a physicist co-chair. Need a Virgo person.
$Id: minutes-2009-02-09.html,v 1.2 2009/02/09 16:50:15 jolien Exp $